In a AP the sum of the three postive consecutive number is 21 and the sum of their sqaure is 155.find the number?
Answer by addingup(3677)

x+x+1+x+1+1= 21
3x+3= 21
3x= 18
x= 6
6+(6+1)+(6+1+1)= 21
6+7+8= 21
21= 21 We have the right answer
--------------------------
x^2+(x+1)^2+(x+2)^2= 155
3x^2+3^2= 155
3x^2= 146
3x= 12.08
x= 4.02
4.02+5.02+6.02= 21
15.06= 21 This is incorrect. Check your numbers, 155 is not the correct number or you're missing critical information.